# Operator entanglement in $\mathrm{SU}(2)$-symmetric dissipative quantum many-body dynamics

### v2

@@ 5.3 Symmetry-resolved operator entanglement in magnetization sectors
As we expected, the probabilities $p_{S_{z}}$ indeed are approximately Gaussian at late times, i.e., $p_{S_{z}}\simeq e^{-S_{z}^{2}/2\delta^{2}}/\sqrt{2\pi\delta^{2}}$; see Fig. 3(a). In Fig. 3(b), we also present the time dependence of the variance $\delta$, which follows the power law $\sim(tJ)^{\alpha}$ at late times and explains the logarithmic growth behavior of $-\sum_{S}p_{S}\log_{2}p_{S}$ shown in Fig. 2(c). The exponent $\alpha$ in general depends on the dissipation strength and for large $\gamma$ is close to the value $0.25$ predicted for the $\mathrm{U}(1)$-symmetric dissipative quantum dynamics with dephasing [38]; see the insert in Fig. 3(b). Obviously, the classical Shannon entropy of the probabilities in magnetization sectors, $-\sum_{S_{z}}p_{S_{z}}\log_{2}p_{S_{z}}\simeq\log_{2}\delta+\log_{2}\sqrt{2\pi e}$, which follows a logarithmical growth law with prefactor $\alpha$, also cannot capture the whole prefactor $\eta$ of the logarithmic growth of operator entanglement at late times; cf. Fig. 1(d).
@@
To identify the behavior of $S_{\mathrm{op},S_{z}}$, we first consider the difference of symmetry-resolved operator entanglement between the finite magnetization sector and $S_{z}=0$ sector by defining $\Delta S_{\mathrm{op},S_{z}}\equiv S_{\mathrm{op},S_{z}}-S_{\mathrm{op},S_{z}=0}$. The results for $\gamma=0.05J$ are presented in Fig. 3(c), where as the time increases $\Delta S_{\mathrm{op},S_{z}}$ decays to very small values. This suggests that the symmetry-resolved operator entanglement $S_{\mathrm{op},S_{z}}$ will take the same value at long times for all of the magnetization sectors. Moreover, we find that $\Delta S_{\mathrm{op},S_{z}}$ scaled by $1/S_{z}^{2}$ collapses almost onto each other at late times for various magnetization sector $S_{z}$ and can be captured by the function $(a+btJ)^{-c}$ with $c\approx 1$; see the insert in Fig. 3(c). As the contributions of the symmetry-resolved operator entanglement can be rewritten as $\sum_{S_{z}}p_{S_{z}}S_{\mathrm{op},S_{z}}=S_{\mathrm{op},S_{z}=0}+\sum_{S_{z}\neq 0}p_{S_{z}}\Delta S_{\mathrm{op},S_{z}}$, where the second term $\sim(tJ)^{2\alpha-c}$ vanishes at long times, this suggests that the late-time behavior of the averaged symmetry-resolved operator entanglement can be captured by $S_{\mathrm{op},S_{z}=0}$, which, just like the total operator entanglement $S_{\mathrm{op}}$, after the initial rise and fall, increases again in a logarithmic manner at late times; see Fig. 3(d). Together with the classical Shannon entropy of the probability distribution $p_{S_{z}}$, this explains the long-time behavior of the total operator entanglement in our $\mathrm{SU}(2)$-symmetric dissipative quantum dynamics.
@@
Since the operator entanglement dynamics in our $\mathrm{SU}(2)$-symmetric case can be fully understood from the corresponding $\mathrm{U}(1)$ subsymmetry, the above results show evidence that the logarithmic growth of operator entanglement at long times is a generic behavior of the dissipative quantum many-body dynamics with $\mathrm{U}(1)$ as the symmetry or subsymmetry. This behavior also holds for more broad dissipations beyond dephasing, which is valid even for the open quantum systems with only $\mathrm{U}(1)$ symmetry, as we demonstrated below by breaking the symmetry of our quantum dynamics to $\mathrm{U}(1)$.
@@ 5.4 Symmetry breaking from SU(2) to U(1)
6 Conclusion
In conclusion, we have studied the far-from-equilibrium dynamics of operator entanglement in a dissipative quantum many-body system with $\mathrm{SU}(2)$ symmetry. We find that after the initial rise and fall, the operator entanglement increases again in a logarithmic manner at late times. This behavior can be fully understood from the corresponding $\mathrm{U}(1)$ subsymmetry by considering the symmetry-resolved operator entanglement, which, unlike the $\mathrm{U}(1)$-symmetric case with dephasing, also contributes to the growth of operator entanglement in addition to the classical Shannon entropy associated with the probabilities for the half system being in different symmetry sectors. Our results show evidence that the logarithmic growth of operator entanglement at late times is a generic behavior of dissipative quantum many-body dynamics with $\mathrm{U}(1)$ as the symmetry or subsymmetry and for more broad dissipations beyond dephasing. In the future, it would be interesting to further test this conjecture by investigating more symmetries like $\mathrm{SU}(N>2)$, $\mathrm{SO}(N)$, and $\mathrm{SP}(2N)$. Moreover, in addition to the strong symmetries considered so far, it would be meaningful to study how the presence of weak symmetries impacts the entanglement dynamics of open quantum systems [62, 63].
Acknowledgements
[…]
@@ Appendix A Details on numerical convergence
We note that in the plots of prefactor $\eta$ and offset $S_{0}$ of $S_{\mathrm{op}}(t\to\infty)=\eta\log_{2}(tJ)+S_{0}$ shown in Figs. 1(c) and 1(d), which are obtained by fitting the data set $\{(\log_{2}(tJ),S_{\mathrm{op}}(t))\}$ for $t=t_{0},t_{0}\pm\delta t$ via a linear function, the results for small $\gamma=0.05J$ and $\gamma=0.10J$ exhibit an artificial wiggle behavior at late times. This arises from numerical errors due to the finite time step and maximal bond dimension. Especially, the later has a more significant influence as the artificial wiggle behavior is highly reduced when the dissipation strength $\gamma$ increases, which reduces the operator entanglement and increases the accuracy of our numerical simulation with fixed maximal bond dimension $\chi=50000$. To remove this artificial wiggle behavior of $\eta$ and $S_{0}$ for these small $\gamma$’s, we could choose a smaller time step and/or a larger maximal bond dimension to further increase the accuracy of our numerical simulations, which, however, requires much more expensive computational resources. Moreover, as we show in the main text, most properties of the operator entanglement, e.g., the late-time logarithmic growth and probability distribution of different symmetry sectors, have already been captured very well in our current numerical simulation. This artificial wiggle behavior at small $\gamma$’s does not affect our results.

### v3

@@ V.3 Symmetry-resolved operator entanglement in magnetization sectors
As we expected, the probabilities $p_{S_{z}}$ indeed are approximately Gaussian at late times, i.e., $p_{S_{z}}\simeq e^{-S_{z}^{2}/2\delta^{2}}/\sqrt{2\pi\delta^{2}}$; see Fig. 3(a). In Fig. 3(b), we also present the time dependence of the variance $\delta$, which follows the power law $\sim(tJ)^{\alpha}$ at late times and explains the logarithmic growth behavior of $\sum_{S}p_{S}\log_{2}p_{S}$ shown in Fig. 2(c). The exponent $\alpha$ in general depends on the dissipation strength and for large $\gamma$ is close to the value $0.25$ predicted for the $\mathrm{U}(1)$-symmetric dissipative quantum dynamics with dephasing [38]; see the insert in Fig. 3(b). Obviously, the classical Shannon entropy of the probabilities in magnetization sectors, $\sum_{S_{z}}p_{S_{z}}\log_{2}p_{S_{z}}\simeq\log_{2}\delta+\log_{2}\sqrt{2\pi e}$, which follows a logarithmical growth law with prefactor $\alpha$, also cannot capture the whole prefactor $\eta$ of the logarithmic growth of operator entanglement at late times; cf. Fig. 1(d).
@@
To identify the behavior of $S_{\mathrm{op},S_{z}}$, we first consider the difference of symmetry-resolved operator entanglement between the finite magnetization sector and $S_{z}=0$ sector by defining $\Delta S_{\mathrm{op},S_{z}}\equiv S_{\mathrm{op},S_{z}}-S_{\mathrm{op},S_{z}=0}$. The results for $\gamma=0.05J$ are presented in Fig. 3(c), where as the time increases $\Delta S_{\mathrm{op},S_{z}}$ decays to very small values. This suggests that the symmetry-resolved operator entanglement $S_{\mathrm{op},S_{z}}$ will take the same value at long times for all of the magnetization sectors. In Appendix B, a proof for this observation is provided, based on the properties of symmetry-resolved operator entanglement $S_{\mathrm{op},S}$ and probability distribution $p_{S}$ in the spin sectors. Although these properties are obtained from numerical simulations, our proof indeed indicates that the late-time same value behavior of $S_{\mathrm{op},S_{z}}$ arises from the $\mathrm{SU}(2)$ symmetry of the dissipative quantum many-body dynamics. Moreover, we find that $\Delta S_{\mathrm{op},S_{z}}$ scaled by $1/S_{z}^{2}$ collapses almost onto each other at late times for various magnetization sector $S_{z}$ and can be captured by the function $(a+btJ)^{-c}$ with $c\approx 1$; see the insert in Fig. 3(c). As the contributions of the symmetry-resolved operator entanglement can be rewritten as $\sum_{S_{z}}p_{S_{z}}S_{\mathrm{op},S_{z}}=S_{\mathrm{op},S_{z}=0}+\sum_{S_{z}\neq 0}p_{S_{z}}\Delta S_{\mathrm{op},S_{z}}$, where the second term $\sim(tJ)^{2\alpha-c}$ vanishes at long times, this suggests that the late-time behavior of the averaged symmetry-resolved operator entanglement can be captured by $S_{\mathrm{op},S_{z}=0}$, which, just like the total operator entanglement $S_{\mathrm{op}}$, after the initial rise and fall, increases again in a logarithmic manner at late times; see Fig. 3(d). Together with the classical Shannon entropy of the probability distribution $p_{S_{z}}$, this explains the long-time behavior of the total operator entanglement in our $\mathrm{SU}(2)$-symmetric dissipative quantum dynamics. We note that in our $\mathrm{SU}(2)$-symmetric case, the symmetry-resolved operator entanglement contributes a logarithmic term with larger prefactor $\eta$ than the classical Shannon entropy of probability distributions (see Fig. 2 and Fig. 3), while in the $\mathrm{U}(1)$-symmetric case with dephasing considered in Ref. [38] the logarithmic growth of Shannon entropy has a larger prefactor (indeed the only nontrivial term).
@@
Since the operator entanglement dynamics in our $\mathrm{SU}(2)$-symmetric case can be fully understood from the corresponding $\mathrm{U}(1)$ subsymmetry, the above results show evidence that the logarithmic growth of operator entanglement at long times is a generic behavior of the dissipative quantum many-body dynamics with $\mathrm{U}(1)$ as the symmetry or subsymmetry, although more analytical or numerical proofs are still required for this conjecture in the future studies. The logarithmic growth behavior of operator entanglement also holds for more broad dissipations beyond dephasing. In the following, by breaking the symmetry of our quantum dynamics to $\mathrm{U}(1)$, we show that this property is valid even for the open quantum systems with only $\mathrm{U}(1)$ symmetry.
V.4 Symmetry breaking from SU(2) to U(1)
@@
We note that it would be interesting to further explore the conditions in which the late-time symmetry-resolved operator entanglement exhibits nontrivial growth or vanishes when the symmetry of dissipative quantum many-body dynamics is broken to $\mathrm{U}(1)$. However, as suggested by Fig. 4, this question highly depends on the initial states and probably also on the terms introduced to break the $\mathrm{SU}(2)$ symmetry of Liouvillian to $\mathrm{U}(1)$. Since we cannot explore all these possibilities in the numerical simulations, we leave this question for future studies.
VI Conclusion
In conclusion, we have studied the far-from-equilibrium dynamics of operator entanglement in a dissipative quantum many-body system with $\mathrm{SU}(2)$ symmetry. We find that after the initial rise and fall, the operator entanglement increases again in a logarithmic manner at late times. This behavior can be fully understood from the corresponding $\mathrm{U}(1)$ subsymmetry by considering the symmetry-resolved operator entanglement. Especially, the probability distribution of different $\mathrm{U}(1)$ sectors also follows the Gaussian distribution observed in the $\mathrm{U}(1)$-symmetric case with dephasing, with the variance growing in a power law. But unlike the latter, in addition to the classical Shannon entropy associated with the probabilities for the half system being in different symmetry sectors, the symmetry-resolved operator entanglement also contributes nontrivially to the late-time growth of operator entanglement in our $\mathrm{SU}(2)$-symmetric case. Our results show numerical evidence that the logarithmic growth of operator entanglement at late times is a generic behavior of dissipative quantum many-body dynamics with $\mathrm{U}(1)$ as the symmetry or subsymmetry and for more broad dissipations beyond dephasing. In the future studies, it would be interesting to further test this conjecture by investigating more symmetries like $\mathrm{SU}(N>2)$, $\mathrm{SO}(N)$, and $\mathrm{SP}(2N)$, and provide more analytical proofs. Moreover, in addition to the strong symmetries considered so far, it would be meaningful to study how the presence of weak symmetries impacts the entanglement dynamics of open quantum systems [62, 63].
Acknowledgements.
We implement the $\mathrm{SU}(2)$ symmetry of the tensors using the package TensorKit.jl [64]. We acknowledge support from: European Research Council AdG NOQIA; MCIN/AEI (PGC2018-0910.13039/501100011033, CEX2019-000910-S/10.13039/501100011033, Plan National FIDEUA PID2019-106901GB-I00, Plan National STAMEENA PID2022-139099NB, I00, project funded by MCIN/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR” (PRTR-C17.I1), FPI); QUANTERA DYNAMITE PCI2022-132919, QuantERA II Programme co-funded by European Union’s Horizon 2020 program under Grant Agreement No 101017733; Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda; MICIU/AEI/10.13039/501100011033 and EU (PCI2025-163167); Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program); Barcelona Supercomputing Center MareNostrum (FI-2023-3-0024); Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, European Commission, European Climate, Infrastructure and Environment Executive Agency (CINEA), or any other granting authority. Neither the European Union nor any granting authority can be held responsible for them (HORIZON-CL4-2022-QUANTUM-02-SGA PASQuanS2.1, 101113690, EU Horizon 2020 FET-OPEN OPTOlogic, Grant No 899794, QU-ATTO, 101168628), EU Horizon Europe Program (This project has received funding from the European Union’s Horizon Europe research and innovation program under Grant Agreement No. 101080086 — NeQST); ICFO Internal “QuantumGaudi” project.
Data availability
The data that support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.
@@ Appendix A Details on numerical convergence
We note that in the plots of prefactor $\eta$ and offset $S_{0}$ of $S_{\mathrm{op}}(t\to\infty)=\eta\log_{2}(tJ)+S_{0}$ shown in Figs. 1(c) and 1(d), which are obtained by fitting the data set $\{(\log_{2}(tJ),S_{\mathrm{op}}(t))\}$ for $t=t_{0},t_{0}\pm\delta t$ via a linear function, the results for small $\gamma=0.05J$ and $\gamma=0.10J$ exhibit an artificial wiggle behavior at late times. This arises from numerical errors due to the finite time step and maximal bond dimension. Especially, the latter has a more significant influence as the artificial wiggle behavior is highly reduced when the dissipation strength $\gamma$ increases, which reduces the operator entanglement and increases the accuracy of our numerical simulation with fixed maximal bond dimension $\chi=50000$. To remove this artificial wiggle behavior of $\eta$ and $S_{0}$ for these small $\gamma$’s, we could choose a smaller time step and/or a larger maximal bond dimension to further increase the accuracy of our numerical simulations, which, however, requires much more expensive computational resources. Moreover, as we show in the main text, most properties of the operator entanglement, e.g., the late-time logarithmic growth and probability distribution of different symmetry sectors, have already been captured very well in our current numerical simulation. This artificial wiggle behavior at small $\gamma$’s does not affect our results.
Appendix B Proof of the late-time same value behavior of $S_{\mathrm{op},S_{z}}$
In this Appendix, we provide a proof for the observation that the symmetry-resolved operator entanglement $S_{\mathrm{op},S_{z}}$ takes the same value at long times for all of the magnetization sectors, based on the properties of symmetry-resolved operator entanglement $S_{\mathrm{op},S}$ and probability distribution $p_{S}$ in the spin sectors that (i) the total operator entanglement $S_{\mathrm{op}}$ grows logarithmically at late times, i.e., $S_{\mathrm{op}}\sim\log_{2}(tJ)$; (ii) the probability distribution $p_{S}$ among different spin sectors follows Eq. (5) with $\delta\sim(tJ)^{\alpha}$ at late times [see Fig. 2(a) and Fig. 3(b)]; (iii) $S_{\mathrm{op},S}$ remains finite at late times [see Fig. 2(b)].
We focus on the region with $S_{z}\geq 0$. The proof for the cases with $S_{z}<0$ is similar. According to the expressions for $S_{\mathrm{op},S}$ ($S_{\mathrm{op},S_{z}}$) and $p_{S}$ ($p_{S_{z}}$) in terms of Schmidt values $\lambda_{S,i_{S}}$ (see Sec. IV), we find the following relations
$p_{S_{z}}$ $=\sum_{S\geq S_{z}}\frac{p_{S}}{2S+1}=\frac{1}{\sqrt{2\pi\delta^{2}}}e^{-S_{z}^{2}/2\delta^{2}},$ (9)
$S_{\mathrm{op},S_{z}}$ $=\frac{1}{p_{S_{z}}}\sum_{S\geq S_{z}}\frac{p_{S}(S_{\mathrm{op},S}-\log_{2}p_{S})}{2S+1}+\log_{2}p_{S_{z}.}$ (10)
We consider the difference bewteen $S_{\mathrm{op},S_{z}}$ and $S_{\mathrm{op},S_{z}-1}$:
$S_{\mathrm{op},S_{z}}-S_{\mathrm{op},S_{z}-1}$ (11)
$=$ $\left(\frac{1}{p_{S_{z}}}-\frac{1}{p_{S_{z}-1}}\right)\sum_{S\geq S_{z}}\frac{p_{S}(S_{\mathrm{op},S}-\log_{2}p_{S})}{2S+1}-\frac{1}{p_{S_{z}-1}}\frac{p_{S=S_{z}-1}(S_{\mathrm{op},S=S_{z}-1}-\log_{2}p_{S=S_{z}-1})}{2S_{z}-1}+\log_{2}\frac{p_{S_{z}}}{p_{S_{z}-1}}$
$=$ $\left(\frac{1}{p_{S_{z}}}-\frac{1}{p_{S_{z}-1}}\right)\sum_{S\geq S_{z}}\frac{p_{S}(S_{\mathrm{op},S}-\log_{2}p_{S})}{2S+1}-\left(1-\frac{p_{S_{z}}}{p_{S_{1}-1}}\right)(S_{\mathrm{op},S=S_{z}-1}-\log_{2}p_{S=S_{z}-1})+\log_{2}\frac{p_{S_{z}}}{p_{S_{z}-1}}$
$\leq$ $\left(\frac{1}{p_{S_{z}}}-\frac{1}{p_{S_{z}-1}}\right)\underbrace{\sum_{S}p_{S}(S_{\mathrm{op},S}-\log_{2}p_{S})}_{S_{\mathrm{op}}}-\left(1-\frac{p_{S_{z}}}{p_{S_{1}-1}}\right)(S_{\mathrm{op},S=S_{z}-1}-\log_{2}p_{S=S_{z}-1})+\log_{2}\frac{p_{S_{z}}}{p_{S_{z}-1}}.$
For the first term, we have
$\left(\frac{1}{p_{S_{Z}}}-\frac{1}{p_{S_{z}-1}}\right)S_{\mathrm{op}}=\sqrt{2\pi\delta^{2}}\left[e^{S_{z}^{2}/2\delta^{2}}-e^{(S_{z}-1)^{2}/2\delta^{2}}\right]S_{\mathrm{op}}\simeq\sqrt{2\pi\delta^{2}}e^{S_{z}^{2}/2\delta^{2}}(2S_{z}-1)S_{\mathrm{op}}/2\delta^{2}\sim\frac{\log_{2}(tJ)}{(tJ)^{\alpha}},$ (12)
which tends to zero at late times. For the second term, on the one hand, $(1-p_{S_{Z}}/p_{S_{z}-1})S_{\mathrm{op},S=S_{z}-1}\simeq(2S_{z}-1)S_{\mathrm{op},S=S_{z}-1}/2\delta^{2}$ vanishes as $t\to\infty$ since $S_{\mathrm{op},S}$ remains finite at late times. On the other hand, we have
$\left(1-p_{S_{z}}/p_{S_{z}-1}\right)\log_{2}p_{S=S_{z}-1}$ (13)
$=$ $\left(1-p_{S_{z}}/p_{S_{z}-1}\right)\log_{2}\left(\frac{2S_{z}-1}{\sqrt{2\pi\delta^{2}}}e^{-S_{z}^{2}/2\delta^{2}}[e^{(2S_{z}-1)/2\delta^{2}}-1]\right)\simeq\frac{2S_{z}-1}{2\delta^{2}}\log_{2}\left[\frac{(2S_{z}-1)^{2}}{\sqrt{8\pi\delta^{6}}}\right]\sim\frac{\log_{2}\delta}{\delta^{2}}\sim\frac{\log_{2}(tJ)}{(tJ)^{2\alpha}},$
which also vanishes at late times. Finally, the third term $\log_{2}(p_{S_{z}}/p_{S_{z}-1})=\log_{2}e^{-(2S_{z}-1)/2\delta^{2}}$ tends to zero at late times. Therefore, we have proved that $S_{\mathrm{op},S_{z}}-S_{\mathrm{op},S_{z}-1}$ vanishes as $t\to\infty$ and $S_{\mathrm{op},S_{z}}$ takes the same value at late times for all $S_{z}$.
We note that although the properties of $S_{\mathrm{op},S}$ and $p_{S}$ are obtained from numerical simulations, the above proofs indeed indicate that the late-time same value behavior of $S_{\mathrm{op},S_{z}}$ arises from the $\mathrm{SU}(2)$ symmetry of the dissipative quantum many-body dynamics.